OpenStudy (anonymous):

plz help explain. What is the solution of log^(2x – 3)125 = 3 ? answers a)x=1/3 b)x=1 c)x=7/3 d)x=4

OpenStudy (anonymous):

Well, here this means such: \[(2x-3)^3 = 125\] Can you proceed from here?

OpenStudy (anonymous):

so it means that I mulitply (2x-3) 3times?

OpenStudy (anonymous):

Yes. Also hint: 125 is a perfect cube.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

It is not 3(2x-3) it is (2x-3)(2x-3)(2x-3)

OpenStudy (anonymous):

once I get the answer for that I can dvide?

OpenStudy (anonymous):

Well, do you know what the perfect cube is for 125? (Do you know what a perfect cube is?)

OpenStudy (anonymous):

5√5?

OpenStudy (anonymous):

1*1*1 = 1 2*2*2 = 8 3*3*3 = 27 4*4*4 = 64 5*5*5 = 125

OpenStudy (anonymous):

okay so it would be 5^3

OpenStudy (anonymous):

Right, so you have: \[(2x-3)^3 = 5^3\] What happens to the cubes?

OpenStudy (anonymous):

you would put (2x-3)^5=3 don't you switch them?

OpenStudy (anonymous):

No, this means "The base 2x-3 logarithm of 125 is 3". So it is \[(2x-3)^3 = 125\]

OpenStudy (anonymous):

so what changed?

OpenStudy (anonymous):

Nothing at all. You've simply rewritten 125 as 5^3 \[(2x-3)^3 = 5^3\] So then what happens to the cubes (the ^3's)

OpenStudy (anonymous):

okay, I understand that. umm don't they become one? srry if I'm not getting where your comming from I just started learning this.

OpenStudy (anonymous):

Its okay. Well, to eliminate a cube, you take the cube root. Similarly to eliminate a square root, you take the square. To eliminate any exponent you can take the root of the exponent. In simplest terms, the cubes simply go away.

OpenStudy (anonymous):

ahh! okay gottcha. so now it becomes (2x-3)=5?

OpenStudy (anonymous):

Thats right, and solve from there.

OpenStudy (anonymous):

answer is 4 thank you very much.

OpenStudy (anonymous):

That's correct. No problem.